Suppose that a monopoly

1) faces the following (normal) demand function: Q = 80 - 2P

2) faces no fixed costs (FC)

3) has constant MC = $10 unit

Answer the following question (note you will need to create your own graph to answer these questions):

What are the profits earned by this monopolist? Note, do not include a dollar sign ($) in your answer.
Answer #1

Find the inverse demand function

P = 80/2-Q/2

P = 40 - 0.5Q

Marginal revenue function same as demand function but with twice the slope

MR = 40 - Q

Marginal cost is $10

At the profit-maximizing level of output marginal revenue and marginal cost should be equal

40 - Q = 10

Q = 30 units

P = 40-0.5*30 = $25

Therefore the price will be $25 and the quantity will be 30 units. There is no fixed cost which means that marginal cost is same as average cost.

Profit = (P-AC)*Q = (25-10)*30

= $450

The profit earned by this monopolist is 450.